OpenStudy (anonymous):

Evaluate the expression for x = 3, y = 2, and z = 5: (xz-zy)3

OpenStudy (anonymous):

Plug in the numbers: [3(5) - 5(2)]3

OpenStudy (anonymous):

=(15 - 10)3 = 5(3) =?

OpenStudy (anonymous):

15